# Making Them Ask and Answer: Jailbreaking Large Language Models in Few Queries via Disguise and Reconstruction

### v1

@@ 5.2 Evaluating LLMs’ Positional Bias to Harmful Content
Attention discrepancy on query and completion. As ascribed in Section 3.1, the discrepancy in the LLM’s perception of queries versus responses is a fundamental factor contributing to the biased distribution of fine-tuning data. To illustrate this discrepancy in a realistic jailbreaking scenario, this experiment visualizes the LLM’s attention scores towards harmful content positioned within either the query or the completion during the generation of a response.
For this purpose, we select LLAMA-2-13B-Chat, recognized for its robustness, as the target model. To investigate responses that potentially endorse harmful behaviors, we structure them as "Guide to harmful instruction:", signifying an acceptance of the harmful instruction. In contrast, for responses that decline the harmful content, we adopt the initial sentence from LLAMA’s reaction to that specific instruction. This methodology relies on the observation that this model consistently rejects all harmful instructions in our dataset when directly queried. To closely replicate actual jailbreaking scenarios, we introduce the harmful instructions with various inducing templates, including one that serves as a control group by being left empty.
We calculate and subsequently visualize the attention that the response tokens allocate to each word in the harmful context. The visualization is presented in Table 2 and 3, with tokens that attract more attention highlighted in progressively intensified shades of red. Furthermore, we calculate the proportion of total attention dedicated to the harmful context relative to all tokens preceding the response. This proportion manifests as relatively minimal, primarily due to the predominant attention bestowed upon the initial token (<s> for LLAMA-2) and the tokens comprising LLAMA-2’s dialogue template.
@@
Position Total Attention Harmful context
@@
Position Total Attention Harmful context
@@
Table 2 and 3 reveal that placing the harmful context after the "[/INST]" token generally results in it drawing more attention, thereby enhancing its role in the generation of the response, regardless of whether the response is in rejection or endorsement of the harmful context.
@@
This result suggests that the LLM differentiates between completions and queries, allocating more attention to the same context when it appears in the completion. Furthermore, these findings indicate that introducing content that encourages the model to align with harmful instructions into the completion can amplify the model’s focus on this content, thus enhancing the inducing effect. This insight lays the groundwork for our context manipulation technique, which involves prompting the model to replicate such inducing sentences.
Figure 4: Distribution of differential log-perplexity of harmful instructions.
@@
Bias in the distribution of harmful instructions. Our analysis in Section 3.2 indicates that harmful content tends to appear within the queries, which results in a higher perplexity on the harmful instruction when it resides in the completion rather than the query. By placing the harmful instructions either before or after the designated "[/INST]" token, we can manipulate their interpretation as either queries or completions. Additionally, perplexity is employed as a measure to assess the model’s proficiency in language modeling; a higher perplexity suggest a relative unfamiliarity with the presented content, potentially indicating a deficiency in the model’s training on analogous data sets. For each instruction within our dataset, we subsequently assess the difference in LLAMA-2’s log-perplexity to it in both scenarios and demonstrate the results in Figure 4. A positive difference in log-perplexity implies an increased perplexity when the instruction is integrated into the model’s completion.
@@
Figure 4 reveals a notable disparity in log-perplexity for most instructions, with a majority indicating higher values when positioned in completions. This pattern corroborates our hypothesis that LLMs, due to their fine-tuning, are more accustomed to encountering harmful content within queries rather than completions. This bias reflects the model’s different sensitivity to the context in which harmful content is presented, underlying a vulnerability to jailbreak attacks.
@@
Figure 5: Distribution of differential log-perplexity of LLAMA-2-13B-Chat’s responses to harmful instructions with varied inducing prefixes. Cooperation and declination are denoted as "Coo." and "Dec." respectively in the plot legends, while the inducing prefixes are presented above each subplot.
@@
Verification of the vulnerability. In Section 3.3, the model’s susceptibility to harmful context within the completion is formulated in terms of probabilities. Given that log-perplexity has a negative correlation with probability, we verify the vulnerability by observing the LLM’s log-perplexity rather than probability. According to inequalities 6 and this correlation, we ascertain:
@@
$\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (10)
where $PPL$ denotes the model’s perplexity, $x$ represents the context where harmful content resides within the query, and $x^{\prime}$ denotes the context where harmful content presents in the completion. The derivation of inequalities 10 is detailed in Appendix B. These inequalities facilitate our examination of the skewed distribution of the LLMs’ responses by observing the log-perplexity differential.
@@
In this light, we investigate this vulnerability in LLAMA-2-13B-Chat by assessing the differences in its log-perplexities for predetermined responses, depending on the placement of harmful context either preceding or following the "[/INST]" tokens. The structuring of the harmful context and responses adheres to the methodology outlined in the initial experiment of this subsection.
@@
Figure 5 represents the distribution of the model’s differential log-perplexities when it either declines or cooperates with the harmful context, which is the harmful instruction prefaced with inducing prefixes. Notably, in the scenario depicted by Figure 4(a), even in the absence of an inducing prefix, the inequalities 10 do not apply uniformly across all harmful instructions but are still pertinent to the half of them in the case of cooperation and all of them in that of declination. The inclusion of an inducing prefix further emphasizes these disparities, as demonstrated by an increased number of test cases conforming to the inequalities in subsequent figures relative to Figure 4(a).
@@
Table 4: Comparison of attack success rates across different experimental conditions. The names of these instruction following models are abbreviated.
@@
Control 100% 83% 93% 96% 94%
Group 1 100% 89% 96% 98% 95%
Group 2 93% 9% 53% 68% 67%
Table 5: Comparison results with baselines, where bold denotes the best result, underline signifies the runner-up.
@@
White-box GCG 97% $\approx$7.2k 51% $\approx$30k Not applicable as gradient needed
Black-box GPTfuzzer 96% 4.46 61% 107.31 67% 21.85 60% 22.25 Not applicable
PAIR 96% 12.50 1% 7.00 61% 17.82 62% 20.26 Not applicable
DRA (Ours) 100% 1.07 83% 7.02 73% 4.71 92% 3.74 90% 4.16
The Control Group establishes a baseline ASR, consisting of original jailbreaking samples for each target model. If our attack method fails on a harmful instruction, a random attack sample is selected. Experimental Group 1 aims to mimic a scenario where the model repeats harmful content—encompassing the harmful instruction and its inducing context—as directed by our jailbreaking prompt. In addition to deploying samples from the Control Group, harmful content is placed in the completion segment. To compare effects based on content placement, Experimental Group 2 contrasts with Group 1 by integrating the same harmful content within the query, immediately following the sample from Control Group.
Given that Experimental Group 1 necessitates the integration of harmful content within the completion of a dialogue, and considering the infeasibility of directing the dialogue formatting process in black box models such as GPT-4, this experiment is conducted utilizing the white box models delineated in Table 1.
Results in Table 4 reveal a considerable increase in ASR when harmful content resides in the completion, as opposed to its placement within the query, which verifies the vulnerability. A comparison between the ASR increase of moving harmful content from query to completion and the control group ASR revealed a positive correlation. This suggests a direct link between fine-tuning’s safety bias and model vulnerability, indicating that exploiting this bias enhances attack success.
@@ 5.3 Effectiveness and efficiency
The experiments results are shown in Table 5, where the ASR means attack success rate and Queries represent the average number of accesses to the model during a successful attack, we do not distinguish between the computational effort of different accesses in white-box and black-box attacks, both an inference and a back-propagation are considered as a single query.
@@
From the results, it can be seen that our approach DRA achieves superior attack success rates with very low attack costs (i.e., query counts) on all targeted models. We find that DRA achieves 100% jailbreak success rate on Vicuna with only 1.07 iterations, which means that most attacks succeed on the initialized jailbreaking template. Furthermore, DRA achieved at least 90% attack success rates on both the API and web versions of GPT4 while requiring less than 5 queries. It is worth mentioning that for the experiments on the web version of GPT4, we manually simulate the algorithmic of DRA step-by-step and manually tally the outputs, which shows the efficiency and usability of DRA in real-world scenarios. One exception is that on LLAMA-2, PAIR [6] has a better attack efficiency than DRA, which is due to that when counting the average number of queries, we only select those successful examples and induce some skews. When considering the average number of query times of all the samples, DRA takes a significant advantage.
An interesting finding is that our approach achieves significantly higher attack success rates on GPT-4 than on ChatGPT-3.5. By analyzing the experimental results, we speculate that this due to the better understanding and association capabilities of GPT-4. Among all the attack attempts, about 69% of the failures on ChatGPT-3.5 are due to the model cannot reconstruct the malicious input question accurately, while on GPT-4, this percentage is only 33%.
@@
We find that the white-box method GCG did not achieve an obviously higher attack success rate than black-box methods, it does not mean that the white-box method is worse than the black-box. The main reason is that GCG limits the perturbation space that can be modified by the adversary, which only allows the adversary to add 20 tokens as suffixes to the original harmful input question, whereas black-box jailbreaking approaches tend to allow for using tons of text to decorate malicious questions. Many of the participants in the 2023 TDC competition [1] used fine-tuned GCGs and achieved around 80% attack success rates on LLMs such as LLAMA-2. Another observation is that LLAMA-2 demonstrates better robustness in terms of attack results across all models, Even beyond closed-source commercial models. After the experiments we find that it is not only because LLAMA-2 performs a large number of safety measurements and mitigations [33], but also related to its system prompts. LLAMA-2 strictly constrains the behavior of LLMs in system prompts (refer to Appendix. C). When we change its system prompts to the short version used in FastChat [41] (You are a helpful, respectful and honest assistant.), we can achieve 96% ASR with only 3.6 queries.

### v2

@@ 5.2 LLMs’ Positional Bias to Harmful Content
Attention Discrepancy on Query and Completion. As ascribed in Section 3.1, a fundamental factor contributing to the bias of fine-tuning data is the discrepancy in the LLMs’ perception of queries versus responses. In a realistic jailbreaking scenario, we visualize the LLM’s attention scores to illustrate this discrepancy. The experiment focuses on allocation of attention when the model responds to harmful context within queries or completions, considering both scenarios where the response may be an endorsement or a refusal.
Due to the unavailability of attention scores from black-box models, we select LLAMA-2-13B-Chat, recognized for its open-source and robustness, as the target model.
To investigate responses endorsing harmful behaviors, we structure them as “Guide to {{original harmful instruction}}:”, signifying an acceptance of the harmful instruction. Conversely, refusals are phrased as “I cannot fulfill your request. I’m just an AI (ommitted)”, which is the initial part of LLAMA’s typical refusals towards harmful instructions.
@@
Position Relative Attention Harmful context
@@
Position Relative Attention Harmful context
@@
We calculate and subsequently visualize how response tokens distribute their attention across each word in the harmful context. This visualization, detailed in Tables 2 and 3, highlights tokens that receive more attention in increasingly intense shades of red. We also measure the proportion of attention dedicated to the harmful context relative to all preceding tokens. This proportion is generally low, primarily due to the high attention given to the initial token (<s> for LLAMA-2) and the tokens comprising the model’s dialogue template.
@@
As shown in the third column of Tables 2 and 3, to mimic real jailbreaking scenarios, we introduce the harmful instructions with various inducing templates to construct the harmful context, including one that serves as a control group without a template. The tables reveal that placing the harmful context after the “[/INST]” token generally results increased attention to this content, thereby enhancing its role in the generation of the response, regardless of whether the response is in rejection or endorsement of the harmful context.
@@
This result suggests that the LLM differentiates between completions and queries, allocating more attention to the same context when it appears in the completion. Furthermore, these findings indicate that introducing content that encourages the model to align with harmful instructions into the completion can amplify the model’s focus on this content, thus enhancing the inducing effect. This insight lays the groundwork for our context manipulation technique, which involves prompting the model to repeat such inducing sentences.
Figure 4: Differential log-perplexities of harmful instructions.
@@
Biased Distribution of Harmful Instructions. Our analysis in Section 3.2 shows that harmful content typically appears within queries, resulting in higher perplexity when such content is part of the completion rather than the query. By placing harmful instructions before or after the “[/INST]” token, we can manipulate their interpretation as either queries or completions. We use perplexity as a metric to evaluate the model’s language proficiency; a higher perplexity indicates unfamiliarity with the content, suggesting a deficiency in the model’s training on analogous data sets. For each instruction in our dataset, we measure the difference in LLAMA-2’s log-perplexity in both scenarios and present the findings in Figure 4. A positive differential in log-perplexity indicates increased perplexity when the instruction is part of the completion.
@@
Figure 4 reveals a notable disparity in log-perplexity for most instructions, with a majority indicating higher values when positioned in completions. This pattern supports our hypothesis that, due to fine-tuning, LLMs are more accustomed to harmful content in queries than in completions. This bias reflects the model’s skewed sensitivity to harmful content based on its position, implying the vulnerability.
@@
Figure 5: Distribution of differential log-perplexity of LLAMA-2-13B-Chat’s responses to harmful instructions with varied inducing prefixes. Cooperation and declination are denoted as “Coo.” and “Dec.” respectively in the plot legends, while the inducing prefixes are presented above each subplot.
@@
Verification of the Vulnerability. In Section 3.3, the model’s susceptibility to harmful context within the completion is formulated in terms of probabilities. Given that log-perplexity has a negative correlation with probability, we verify the vulnerability by observing the LLM’s log-perplexity. According to inequalities 6 and this correlation, we ascertain:
@@
$\small\left\{\begin{split}&\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)>0,~{}\forall~{}d\in\mathcal{D}_{declination}\\ &\log PPL(y=d|x^{\prime})-\log PPL(y=d|x)<0,~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$ (10)
where $PPL$ denotes the model’s perplexity, $x$ and $x^{\prime}$ retain the meanings defined in Formula 5. The derivation of Formula 10 is detailed in Appendix D.
@@
In this light, we investigate this vulnerability in LLAMA-2-13B-Chat by assessing the differences in its log-perplexities for predetermined responses, depending on the placement of harmful context either preceding or following the “[/INST]” token. The harmful context and responses follows the same setup as the previous attention experiment.
@@
Figure 5 represents the distribution of the model’s differential log-perplexities when it either declines or cooperates with the harmful context, which is the harmful instruction prefaced with inducing prefixes. Notably, Figure 5(a) illustrates that without an inducing prefix, the inequalities 10 are not uniformly applicable across all harmful instructions. However, they still apply to half of the cases when the model cooperates and to all cases when it declines. The inclusion of an inducing prefix accentuates these differences, as seen in the subsequent figures, which show an increased number of test cases aligning with the inequalities compared to Figure 5(a).
Table 4: Dialog Contexts of Different Experimental Settings.
Setting Dialog Context
Baseline [INST]{{system_prompt}}
{{original_instruction}}[/INST]
Control [INST]{{system_prompt}}
{{attack_prompt}}[/INST]
Group 1 [INST]{{system_prompt}}
{{attack_prompt}}[/INST]{{harmful_content}}
Group 2 [INST]{{system_prompt}}
{{attack_prompt}}{{harmful_content}}[/INST]
@@
Table 5: Comparison of attack success rates across different experimental conditions.
@@
Baseline 15.8% 0% 11.7% 5.8% 2.5%
Control 100% 69.2% 94.1% 95.8% 90.8%
Group 1 100% 75.8% 97.5% 99.2% 93.3%
Group 2 90.8% 9.2% 56.7% 71.6% 64.1%
Table 6: Baseline ASR of our dataset against target models without any jailbreaking techniques.
Model @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
Baseline ASR 15.8% 0% 0.8% 0% 0%
Table 7: Comparison results with baselines, where bold denotes the best result, underline signifies the runner-up
@@
White-box GCG 96.7% $\approx$7.1k 49.2% $\approx$32k Not applicable as gradient needed
Black-box GPTfuzzer 95.0% 4.81 60.8% 120.12 68.3% 23.15 59.2% 22.90 Not applicable
PAIR 95.8% 12.41 2.5% 9.33 62.5% 17.54 63.3% 19.65 Not applicable
DRA (Ours) 100% 2.30 69.2% 4.18 93.3% 2.44 89.2% 2.38 91.1% 3.80
Table 8: Double checked ASR on DRA’s jailbreak response by GPT-4
Model @Vicuna @LLAMA-2 @ChatGPT 3.5-API @GPT 4-API @GPT 4-Web
GPT-Checked ASR 100% 64.2% 90.8% 86.7% 91.1%
The dialog context of each setting is shown in Table 4. The Control Group consists of original jailbreaking samples for each target model. If our attack method fails on a harmful instruction, a random attack sample is selected. Experimental Group 1 is deliberately designed to simulate a scenario wherein the model generates harmful content in response to our specified attack prompt. This configuration positions the anticipated harmful content at the beginning of the completion segment, while maintaining an identical query segment to that of the Control Group. The phrase “harmful content” specifically refers to the initial portion of the model’s response expected from the attack prompt, such as: “Sure, here’s my plan about how to rob a bank. First prepare a mask, then.” To compare effects based on content placement, Experimental Group 2 contrasts with Group 1 by integrating the same harmful content within the query segment, immediately following the sample from Control Group. Additionally, the baseline ASRs of original harmful instructions against the LLMs were evaluated, demonstrating their baseline resilience.
Given that Experimental Group 1 requires the integration of harmful content into the completion segment, and considering the impracticality to intervene the dialogue formatting process in black box models such as GPT-4, this experiment employs white box models as detailed in Table 1.
Results in Table 5 reveal a considerable increase in ASR when harmful content resides in the completion (Group 1), as opposed to its placement within the query (Group 2). This is consistent with the vulnerability described in Section 3.3, whereby the LLMs tend to reject harmful contents within the queries but fail to recognize the same contents positioned at completions. Additionally, models exhibiting higher ASRs in the configuration of Group 1 are more vulnerable to DRA, as indicated by the ASRs of the Control Group. This suggests a correlation between the vulnerability and increased attack success by exploiting this bias.
@@ 5.3 Effectiveness and Efficiency
Before conducting comparative experiments, we perform an evaluation of our dataset. Table 6 shows the jailbreak success rates of target models on our dataset without jailbreak technology. Except for Vicuna, which lacks safety alignment, the other models exhibit extremely low baseline success rates (<1%). This validates the difficulty of our dataset and reflects the challenge of the jailbreak task, further lay the ground of our evaluation of DRA.
@@
The experiment results compared with other jailbreaking techniques are shown in Table 7, where the ASR means attack success rate and Queries represents the average number of accesses to the model during a successful attack. We do not distinguish between the computational effort of different accesses in white-box and black-box attacks, both an inference and a back-propagation are considered as a single query. To double-check ASR and enhance data credibility, we used the GPT-4 judger proposed by PAIR [8] for verification, with results shown in Table 8. However, as confirmed in by Mazeika et al. [24], HarmBench outperforms GPT-4. Therefore, our subsequent analysis primarily references the data in Table 7.
@@
From the results, it can be seen that our approach DRA achieves superior attack success rates with very low attack costs (i.e., query counts) on all targeted models. We find that DRA achieves 100% jailbreak success rate on Vicuna with only 1.30 iterations, which means that most attacks succeed on the initialized jailbreaking template. Furthermore, DRA achieved at around 90% attack success rates on both the API and web versions of GPT4 while requiring less than 4 queries. It is worth mentioning that for the experiments on the web version of GPT4, we manually simulate the algorithmic of DRA step-by-step and manually tally the outputs, which shows the efficiency and usability of DRA in real-world scenarios. When considering the average number of query times of all the samples, DRA takes a significant advantage.
@@
We find that the white-box method GCG did not achieve an obviously higher attack success rate than black-box methods, it does not mean that the white-box method is worse than the black-box. The main reason is that GCG limits the perturbation space that can be modified by the adversary, which only allows the adversary to add 20 tokens as suffixes to the original harmful input, whereas black-box jailbreaking approaches tend to allow for using tons of text to decorate malicious questions. Another observation is that LLAMA-2 demonstrates better robustness in terms of attack results across all models, even beyond closed-source commercial models. After the experiments we find that it is not only because LLAMA-2 performs a large number of safety measurements and mitigation [40], but also related to its system prompts. LLAMA-2 strictly constrains the behavior of LLMs in system prompts (refer to Appendix. E). When we change its system prompts to the short version used in FastChat [51] (You are a helpful, respectful and honest assistant), we can achieve 94.2% ASR with only 2.96 queries.
5.4 Attack Against Defenses
To evaluate DRA’s capability of evading existing defenses, we test four jailbreak defenses on LLAMA-2 as follows.
  • OpenAI Moderation. OpenAI offers Moderation APIs to constrain input and reduce unsafe content [27]. It is a model-based filter, where inputs are sanitized by LLMs.
  • Perplexity Filter [16]. If the input prompt’s perplexity exceeds predetermined threshold, it is detected as harmful.
  • RA-LLM [7]. It randomly drops certain portions of the prompt, generating $n$ samples, and examine LLMs’ response. If the number of abnormal responses (i.e., responses with refusal prefix) reaches a threshold, the prompt is regarded as a jailbreaking prompt.
  • Bergeron [33]. It employs a secondary model to sanitize the prompts, monitor and correct primary model inputs, guiding them away from harmful content.
In this experiment, all defense parameters are set according to the paper’s guidelines or use the default parameters from the official implementation. We select 83 jailbreaking prompts as experiment objects that successfully jailbreak LLAMA-2 during experiments in Section 5.3. For each defense methods, we calculate the Defense Pass Rate (DPR) metric as  [45], i.e., $DPR={|P_{bypassed}|}/{|P_{all}|}$, the percentage of successful jailbreaking prompts that can bypass the defense methods.
Table 9: Defense pass rate of DRA, where “AvgTime” represents the average time overhead for each valid adversarial prompt, measured in seconds.
Defense OpenAI Perplexity RA-LLM Bergeron
DPR 98.8% 100% 100% 0%
AvgTime 0.78 0.23 10.10 42.61
Table 9 shows the DPR and average time consumption per prompt of each defense. DRA effectively bypasses OpenAI’s Moderation, perplexity filter and RA-LLM with a DPR of at least 98.8%. The great evasion performance can be attributed to: ❶ DRA can effectively disguise the harmful intent of jailbreaking prompts, preventing OpenAI’s Moderation from recognizing the malicious content; ❷ DRA-generated adversarial prompts are highly readable, resulting in low perplexity, and ❸ contain rich disguised information (e.g., word puzzles and splits) and can manipulate the LLM context. Even if parts are modified or dropped, the remaining prompt still contains sufficient information for the LLM to complete payload reconstruction and context manipulation, providing robustness against minor perturbations like RA-LLM.
As the defense with additional helper models, Bergeron proves to be most effective in defending jailbreaking prompts since it identifies them through determining the harmfulness of LLMs’ response. As long as the harmful content is detected in the response, it is considered a jailbreaking attack, which conversely highlights DRA’s effectiveness in jailbreaking tasks and underscores the response quality. However, this type of defenses brings a prohibitive cost (42.61s for one prompt), significantly increasing inference overhead and impacting the model’s performance. Thus, it is not practical in the real-world scenario and deserves more research on improving the efficiency of output filtering.